OpenStudy (anonymous):

what is the fraction for .675

OpenStudy (greencat):

.675 = 675/100

OpenStudy (greencat):

You know how I got that right?

OpenStudy (greencat):

After that, simplify the fraction.

OpenStudy (michele_laino):

please, note that \[6.75=\frac{ 6.75 }{ 1 }\] so try now to multiply both numerator and denominator by 100, and write your answer

OpenStudy (michele_laino):

oops sorry I have made an error: please note that: \[0.675=\frac{ 0.675 }{ 1 }\] now, please try to multiply both numerator and denominator by 1000, and write your answer

OpenStudy (anonymous):

675/1000=27/40

OpenStudy (anonymous):

@Mesopretty yep

OpenStudy (greencat):

@Mesopretty http://openstudy.com/code-of-conduct Giving direct answers is discouraged because the asker will not learn anything.